OpenStudy (xapproachesinfinity):

prove that \[\int \frac{dx}{(x^2+1)^{n+1}}=\frac{x}{2n(x^2+1)^n}+(1-\frac{1}{2n})\int \frac{dx}{(x^2+1)^n}\]

Nnesha (nnesha):

proved by @nnesha

OpenStudy (adamk):

proved by @AdamK

OpenStudy (xapproachesinfinity):

this supposed to be by parts i tried dozen of things hahha

OpenStudy (xapproachesinfinity):

@Nnesha you want to try haha

OpenStudy (xapproachesinfinity):

@freckles another problem here lol

OpenStudy (xapproachesinfinity):

hehe just give it a try you never know lol

OpenStudy (xapproachesinfinity):

hahah :)

OpenStudy (freckles):

\[\int\limits_{}^{}\frac{x}{(x^2+1)^{n+1}} dx \\ \frac{1}{2} \int\limits \frac{ du}{u^{n+1}}=\frac{1}{2}\frac{u^{-n-1+1}}{-n-1+1}+C= \frac{-1}{2n u^n}+C=\frac{-1}{2n(x^2+1)^{n}} +C \\ (\frac{1}{x})'=\frac{-1}{x^2} \\ \int\limits \frac{1}{x} \cdot \frac{x}{(x^2+1)^{n+1}} dx= \\ \frac{1}{x} \cdot \frac{-1}{2n (x^2+1)^n}- \int\limits (\frac{-1}{x^2}) \frac{-1}{2n(x^2+1)^n} dx \\ \] hmmm but this doesn't looks like it helps

OpenStudy (xapproachesinfinity):

we don't have x on the top for that integrand it is just \[\int \frac{dx}{(x^2+1)^{n+1}}\]

OpenStudy (xapproachesinfinity):

@perl some help :)

OpenStudy (freckles):

Yeah I know I multiplied by x/x

OpenStudy (freckles):

see bottom 2 lines

OpenStudy (xapproachesinfinity):

oh you got me for the first line haha i thought you did something different actually i tried that me and some friends but we didn't really got an interesting thing that look like what we needed

OpenStudy (freckles):

I don't know I didn't get an x in the numerator :(

OpenStudy (freckles):

for the first term thing

OpenStudy (xapproachesinfinity):

that x and x^2 make a problem for us

OpenStudy (xapproachesinfinity):

i tried with u=x^2+1 then use by parts later

OpenStudy (xapproachesinfinity):

my integral \[\int \frac{du}{\sqrt{u-1} u^{n+1}}\]

OpenStudy (xapproachesinfinity):

of times 1/2

OpenStudy (xapproachesinfinity):

forgot 1/2 there

OpenStudy (anonymous):

The proof is attached. @xapproachesinfinity @freckles @ganeshie8

OpenStudy (xapproachesinfinity):

thanks a lot @eliassaab

OpenStudy (anonymous):

YW